Can you help me with these properties? I want to say one of them is communative. What Property is demonstrated by the following statements?
1. uv(w + x) = uvw + uvx
2. a + cb = a + bc

Answer by Earlsdon(6103)

1. uv(w + x) = uvw + uvx This is an example of the distributive property of multiplication over addition.
2. a + cb = a + bc This is an example of the commutative (note the spelling) property.